1. Query Processing (30 pts): A file of 4096 blocks is to be sorted with an available buffer space of 5 blocks. Assume that reading a block has 1 unit cost and writing a block has 1 unit cost.

What is the minimal cost of the sort-merge algorithm in units?
Answer #1

Let n be number of runs in merge phase

Given, number of blocks in file, b = 4096

Available buffer size, a = 5

now, n = b/a = 4096/5 = 819

Degree of Merging = x = Min(n,a-1) = Min(819,4) = 4

Number of passes in merge phase = Logxn = Log4819 = 4.84

therefore cost of sort-merge algorithm in units = 4.84*2*4096 = 39649.28